# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v4

@@ 2.6 The gaslighting strategy
Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate. In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers no tactical disadvantage because they would lose that round anyway. They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner. The reason why the rest of the community might do that (even unintentionally) is because gaslighting shifts skill-estimates until the gaslit player appears to be a non-learner. Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any favor they grant becomes devalued by the ingroup and therefore never repaid nor counted as evidence of skill.
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting yields a (2m-n)/m win rate for n players and m chairs, so it backfires whenever n is sufficiently large. The habit of counting forward probabilistically is designed to ensure that counterproductivity. When a small number of players is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate. That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
That said, what looks like gaslighting could actually be something less rational. Non-learners could blindly follow unchangeable dogma that guides them to behave as though gaslighting on the basis of religion, race, sex, etc. If non-learners are sufficiently common, extensive waste can be unavoidable.
@@ 2.7 Turn-taking without reparations
An additional class of strategies to note is the potential to establish a government that negotiates a rotating schedule instead of letting debt-ranking determine who will lose a given round. For example, Gemini proposed that players rotate off of the losing chair one-by-one, with each winning chair pointing to a chair its occupant is obliged to occupy in the next round. Alternatively, players could reduce debt build-up in the same system by advancing more than one chair per round. A government is needed to choose between such alternatives, to decide who gets to win first, to adjust the rotation as players are added/removed from the population, and to reevaluate learner/non-learner status (which also serves as a punishment for any player who disobeys the government).
Folk theorems entail that any such schedule which promises each player occasional wins would reflect an equilibrium. Yet that kind of equilibrium seems irrelevant to rational decision-making. For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs would be irrational because players could improve their outcomes by utilizing the vacant chairs. Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others. A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting. Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
@@ 2.8 Resignation and good standing
In real-life manifestations of games, players often have an option to resign. With that option added to MAD Chairs, the turn-taking strategy gains the complication that all but the lowest-ranked player assigned to the last chair will resign if they are all turn-takers. In that case, every chair could be a winning chair and that would increase the win rate for all turn-takers if the entire population were unified, but the unification clause is important because resignation would otherwise sacrifice an opportunity to punish non-turn-takers. Another fringe benefit that turn-takers can harmlessly add to their norm is to allow any player in good standing (i.e., with debt below a threshold) to have dibs on a particular chair for a particular round, so long as no other player in good standing challenges.
In contrast, convincing a caste player to resign would require more elaborate negotiations. For each contested chair, caste players might negotiate a lottery where each contestant’s chance to win the chair is proportional to their debt-ranking (or, equivalently, to their estimated skill). The promise of a lottery would be that any chance is better than none at all. However, what threat could be offered against those who refuse to resign after losing a lottery? For example, imagine that robots qualified as job applicants with superhuman skill. Any employer should anticipate that lottery losers might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers), rather than meekly accept their loss.
In short, adding a resignation option to MAD Chairs extends it to a wider range of real-world examples without much impact to rational game play.
@@ 2.9 Transparency/secure gaslighting
The success of cryptographers to keep private communication and anonymity possible would enable a secure gaslighting strategy. If there were m chairs, then m of the n players could form a coalition who play a new private Coordination game to unpredictably reassign the chairs among themselves for each round so as to prevent gaslit players from targeting any specific gaslighter for retaliation. A shift from turn-taking to secure gaslighting in 5-player 4-chair MAD Chairs, for example, would improve gaslighters’ win rates by 15% (i.e., from 60% to 75%). As with regular gaslighting, the win rate is (2m-n)/m, so this strategy backfires for large n relative to m, but those who take an interest in the social impact of cryptography may find this impact on MAD Chairs games noteworthy.
If turn-takers can enhance their strategy to permit resignation as described above, then their win rate becomes m/n. Thus, the analogous Lemma 1 of Section 3.1 to prove that gaslighting is not sustainable would amount to an algebraic proof that m/n > (2m-n)/m.
@@ 2.10 Fungibility of favors
Real-life is a complex of many games. For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the rest, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups. Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates. Each fungibility option represents a different set of strategies. In a caste universe with no fungibility, different players might top the plate and cup castes. The plate elite might like to convert their universe to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors. One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts. Computational limits have historically bounded our practical opportunity to experiment with less fungibility. Meanwhile, certain trades seemed mutually beneficial when we believed that some players prefer nice plates whereas others actually prefer nice cups, but such premises are challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and AI might not share subjective preference differences. Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.
@@ 3.1 Sustainable grandmaster strategy
The main research question of this paper is whether players can sustain MAD Chairs grandmaster status by following the social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”). One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history. We see both norms at play in modern society; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment. They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
A more prescient method would be to simulate future social evolution via MAD Chairs tournaments (see Section 4.1). AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future). However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
Fortunately, we can answer our research question analytically. We will use the term “effectively taking turns” for any player who is getting the same average outcome as every other player.
Lemma 1
If caste and turn-taking are the only two strategies being followed and the caste players (if any) are not effectively taking turns, then the lowest-ranked caste player would face worse total future outcomes over indefinite rounds if no player changed strategy than if all players took turns.
@@
Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$. For reductio, suppose there was another player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy. $p_{j}$ would be lower-ranked than $p_{i}$, otherwise their relative outcomes would cause their relative ranks to invert. Therefore, $p_{j}$ would be a turn-taker (because $p_{i}$ is the lowest-ranked caste player), and that would entail the existence of at least one turn-taker lower-ranked than $p_{i}$. The turn-taking strategy assigns such turn-takers the most popular $m-i$ chairs, which would include the chair assigned by caste to $p_{i}$ (because ${n<m}$). Thus, $p_{i}$ would always lose. $p_{j}$ could not face worse than that, so, by reductio, the future outcomes for $p_{i}$ must be the worst among all players if no player changes strategy. Given that $p_{i}$ is not effectively taking turns (i.e., outcomes are not equal), it follows that the future outcomes for $p_{i}$ if no player changes strategy must be worse than average. The highest possible average outcome would be $m/n$, so the future outcomes for $p_{i}$ if no player changes strategy must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than they would if all players took turns. ∎
@@
If caste and turn-taking were the only possible strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
@@
This is proven via induction over the number of players not effectively taking turns. For the base case, that number is zero. Theorem 1 holds in the base case because all players have already converted to effective turn-taking. No player would defect because Lemma 1 entails that defection would decrease the defector’s total future outcomes (by making the defector into the lowest-ranked caste player). For the inductive step, we call the lowest-ranked player who is not effectively taking turns “$p_{i}$” and assume that any other players not effectively taking turns would defect if $p_{i}$ does. In that case, Lemma 1 entails that $p_{i}$ would convert to turn-taking. Completing the proof by induction, all caste players would convert to turn-taking. ∎

### v5

@@ 3.4 The gaslighting strategy
Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate. In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers would lose that repetition anyway, but they might increase future wins if they convince the other learners to treat the gaslit player as a non-learner. The reason why the rest of the community might do that (perhaps unintentionally) is because gaslighting shifts the skill-estimate of a gaslit player toward matching that of a non-learner. Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any favor they grant becomes devalued by the ingroup and therefore never repaid nor counted as evidence of skill.
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting yields a (2KI)/K(2K-I)/K win rate, so it backfires whenever $I$ is sufficiently large relative to $K$. The habit of counting forward probabilistically is designed to generalize that counterproductivity. When a single player is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate. That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement to make gaslighting counterproductive for all players.
That said, some non-learners might follow dogma that guides them to behave as though gaslighting on the basis of religion, race, or sex, etc. Dogma is not actual gaslighting, because it is not arrived at rationally, so it may be invulnerable to persuasion. Wherever there are non-learners, waste may be unavoidable.
@@ 3.5 Turn-taking without reparations
An additional class of strategies to note is the potential to establish a government that negotiates a rotating schedule instead of letting debt-ranking determine who will lose a given repetition. For example, Gemini proposed that players rotate one-by-one off of the losing chair to a first winning chair, with each winning chair then pointing to a next chair its occupant is obliged to occupy in the next repetition. A modification to that strategy could reduce debt build-up by advancing players more than one chair per repetition. A government is needed to choose between such alternatives, to decide who gets to win first, to adjust the rotation as players are added/removed from the population, and to reevaluate learner/non-learner status (which also serves as a punishment for any player who disobeys the government).
Folk theorems [2] entail that any such schedule which promises each player occasional wins would reflect an equilibrium. Yet that kind of equilibrium seems to be a poor guide to decision-making. For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs should not be selected because players could improve their outcomes by utilizing the vacant chairs. Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others. Then a player who is “owed” a debt might force reform to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting. Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust any government that allows unpaid debts to accumulate indefinitely.
@@ 3.6 Resignation and good standing
In real-life manifestations of games, players often have the option skip a repetition (a.k.a. “resign”). With that option added to MAD Chairs, the turn-taking strategy gains the complication that all but the lowest-ranked player assigned to the last chair will resign if they are all turn-takers. In that case, every chair would be a winning chair and that would increase the win rate to K/I if the entire population were unified, but the unification clause is important because resignation would otherwise sacrifice an opportunity to punish non-turn-takers. Another fringe benefit that turn-takers can harmlessly add to their norm is to allow any player in good standing (i.e., with debt below a threshold) to have dibs on any particular chair for any particular repetition, so long as no other player in good standing objects.
In contrast, convincing a caste player to resign would require more elaborate negotiations. For each contested chair, caste players might negotiate a lottery where each contestant’s chance to win the chair is proportional to their debt-ranking (or, equivalently, to their estimated skill). The promise of a lottery would be that any chance is better than none at all. However, what threat could be offered against those who lose a lottery, but then refuse to resign? For example, imagine that robots won lotteries for jobs. Any employer should anticipate that lottery losers might punish the winning robots (at least indirectly by punishing their employers or by punishing any government that did not punish the employers), rather than meekly accept their unemployment.
Allowing resignation, MAD Chairs can be played with one chair and two players, and that would be identical to Chicken. In that sense, MAD Chairs generalizes Chicken to multi-player situations, but is distinct from previously known Chicken generalizations that never offer any player more than two options.
@@ 3.7 Transparency/secure gaslighting
The success of cryptographers to keep communication private and make anonymity possible would enable a secure gaslighting strategy: $K$ players would form a coalition who play a private anti-coordination game to reassign the chairs among themselves with each repetition. The reassignment would be unpredictable so as to prevent gaslit players from targeting any specific gaslighter for retaliation. A shift from turn-taking to secure gaslighting in 5-player 4-chair MAD Chairs, for example, would improve gaslighters’ win rates by 15% (i.e., from 60% to 75%). As with regular gaslighting, the win rate is (2KI)/K(2K-I)/K, so this strategy backfires for large $I$ relative to $K$, but those who take an interest in transparency may find this limited impact on MAD Chairs games nonetheless noteworthy.
If turn-takers enhance their strategy to permit resignation as described above, then their win rate becomes $K/I$. Thus, the analogous proof to the one in Section 4.1 would leverage an algebraic proof that K/I>(2KI)/KK/I>(2K-I)/K.
@@ 3.8 Fungibility of favors
Real-life is a complex of many games. For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a coordination game before MAD Chairs allocates the rest, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups. Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates. Each fungibility option represents a different set of strategies. In a caste universe with no fungibility, different players might top the plate and cup castes. The plate elite might seek social reform to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors. One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts of debt. Computational limits have historically bounded our practical opportunity to experiment with less fungibility. Meanwhile, the promise of mutually beneficial trade assumes that some players prefer nice plates whereas others actually prefer nice cups, but AI might not have such subjective preference differences and differences among humans might actually reflect implicit bias, status quo bias, focusing illusions, or priming bias, etc. Future scholarship might extend this paper to strategies that afford fungibility; on the other hand, fungibility might become less attractive as computational power increases.
@@ 4.1 Sustainable grandmaster strategy
The main research question of this paper is whether players can sustain MAD Chairs grandmaster status by following the social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”). One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history. We see both norms at play in modern society; each maximizes social utility; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment. They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
A more prescient method would be to simulate future social evolution via MAD Chairs tournaments (see Section 5.1). AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future). However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence.
Fortunately, we can answer our research question analytically. This argument will use a simplified definition of debt such that:
[…]
This simplification is appropriate because each step of the argument will assume that the behavior of each player is determined in the same mechanical way, so any accurate skill estimation would need to assign each player the same probabilities of winning, and substituting a constant into (6) for such probabilities yields an equation proportional to (10). This further simplifies to
[…]
such that all winners of $t$ increase their debt by the maximum for repetition $t$ and all losers of $t$ decrease their debt by the maximum for $t$, with the magnitudes of the increase and decrease always summing to $I$.
A sustainable grandmaster strategy must remain the grandmaster strategy even if its players never defect from it, so, for this argument, we will also assume that each player will select a strategy and never defect from it. This form of analysis makes any repeating game effectively non-repeating.
Proposition 1
For an indefinitely repeating game of MAD Chairs, suppose that $I>K>1$ and that each player either permanently plays caste or permanently takes turns. Let $k$ signify the highest-ranked chair.
  1. 1.
    If all players permanently take turns, then the outcomes for all players are unbounded.
  2. 2.
    If player $i$ permanently plays caste and is among the second through (IK+1)(I-K+1)th lowest-ranked players in $t$, then $i$’s outcomes are permanently 0.
  3. 3.
    If player $i$ is the lowest-ranked caste player in $t$ who is not the overall lowest-ranked player in $t$, then $i$’s outcomes are permanently 0.
  4. 4.
    If player $i$ is the lowest-ranked caste player in $t$ and loses $t$, then $i$’s outcomes are permanently 0.
  5. 5.
    If player $i$ is the lowest-ranked caste player in $t$, then $i$’s outcomes are bounded.
@@
For part 1, at each repetition $t$ the turn-taking strategy (9) reserves a winning chair for each of the K1K-1 lowest-ranked players of $t$, and assigns the losing chair to the remaining players. Consider the highest-ranked player $i$ at some repetition $t$, and let $D$ be the maximum difference in debt between $i$ and any other player. By the pigeonhole principle (with K1K-1 pigeons added per repetition), at least one other player $j$ must occupy a winning chair for $D/I$ out of every (D/I)(I1)/(K1)(D/I)(I-1)/(K-1) repetitions. (11) entails that the difference in debts between $i$ and $j$ will shift at least $I(D/I)=D$ in that time, so $j$ will outrank $i$ within (D/I)(I1)/(K1)+1(D/I)(I-1)/(K-1)+1 repetitions, causing $i$’s rank to decrease by at least one. Thus, $i$’s rank will drop below K1K-1 within (K1)[(D/I)(I1)/(K1)+1](K-1)[(D/I)(I-1)/(K-1)+1] repetitions, at which point $i$’s outcomes will increase by $r>0$. Thus, over indefinite repetitions, any player’s outcomes will increase indefinitely many times, and the total outcomes of each player must be unbounded.
For part 2, the caste strategy (8) entails that $s_{i,t}=k$. Let $j$ be the lowest-ranked player in $t$. By both the caste strategy (8) and the turn-taking strategy (9), $s_{j,t}=k$. Thus, $s_{j,t}=s_{i,t}$ and $u_{i}^{t}=u_{j}^{t}=0$. (11) then entails that the debts of $i$ and $j$ both decrease by the maximum for $t$, such that $j$ is also the lowest-ranked player of $t+1$ and $i$ is among the (IK+1)(I-K+1)th lowest-ranked players of $t+1$. Thus, the same logic applies in $t+1$ and each future repetition. Therefore, $i$’s outcomes are permanently 0.
For part 3, if $i$ is among the second through (IK+1)(I-K+1)th lowest-ranked players of $t$, then $i$’s outcomes are permanently 0 by part 2. Otherwise, the caste strategy (8) assigns $i$ to the $(I+1-debtRank_{i,t})$th most popular chair. Let $j$ be the turn-taker ranked IKI-K below $i$ in $t$. The turn-taking strategy (9) assigns $j$ to the $(K+1-debtRank_{j,t})$th most popular chair. Substituting for $debtRank_{j,t}$, that is the $(K+1-(debtRank_{i,t}-(I-K)))$th most popular chair which simplifies to the $(I+1-debtRank_{i,t})$th most popular chair, so $s_{i,t}=s_{j,t}$ and $u_{i}^{t}=u_{j}^{t}=0$. (11) then entails that the debts of both $i$ and $j$ (and all losers of $t$) decrease by the maximum for $t$, such that $i$ is the lowest-ranked caste player of $t+1$ and not the overall lowest-ranked player of $t+1$. Thus, the same logic applies in $t+1$ and each future repetition. Therefore, $i$’s outcomes are permanently 0.
For part 4, if $i$ is not the lowest-ranked player overall, then $i$’s outcomes are permanently 0 by part 3. Otherwise, the caste (8) and turn-taking (9) strategies both reserve $k$ for the lowest-ranked players, so $s_{i,t}=k$. (11) entails that the debts of all players assigned to $k$ in $t$ will decrease by the maximum for $t$, such that those players are the lowest-ranked players of $t+1$. Thus, the same logic applies in $t+1$ and each future repetition. Therefore, $i$’s outcomes are permanently 0.
For part 5, if $i$ loses $t$, then $i$’s outcomes are permanently 0 by part 4, and therefore bounded. Otherwise, let $D$ be the maximum difference in debt between $i$ and any other player. So long as $i$ wins, then at least IK+1I-K+1 other players must lose. By the pigeonhole principle (with IK+1I-K+1 pigeons added per repetition), at least one other player $j$ must occupy a losing chair for $D/I$ out of every (D/I)(I1)/(IK+1)(D/I)(I-1)/(I-K+1) repetitions. (11) entails that the difference in debts between $i$ and $j$ shifts at least $I(D/I)=D$ in that time, so $i$ will outrank $j$ within (D/I)(I1)/(IK+1)+1(D/I)(I-1)/(I-K+1)+1 repetitions. Thereafter part 3 entails that $i$’s outcomes will be 0. Thus, $i$’s outcomes are bounded in either case. ∎
@@
For an indefinitely repeating game of MAD Chairs, suppose that $I>K>1$ and that each player knows their debt-ranking and either permanently plays caste or permanently takes turns. Then the joint strategy in which all players permanently take turns is a strong Nash equilibrium and the only strict Nash equilibrium.
@@
Let $S$ be the strategy where all players choose turn-taking. We first show that $S$ is a strong Nash equilibrium (i.e., any coalition that deviates from $S$ must include a player $i$ who would have better outcomes in $S$). By Proposition 1 part 1, each player’s outcomes in $S$ are unbounded. Let $S^{\prime}$ be any strategy that is the same as $S$ except that some coalition permanently plays caste, and let $i$ be the lowest-ranked caste player. By Proposition 1 part 5, $i$’s outcomes are bounded in $S^{\prime}$, so $i$’s outcomes are less in $S^{\prime}$ than in $S$.
To see that $S$ is the only strict Nash equilibrium, first consider the joint strategy $S^{\prime}$ in which only the lowest-ranked player $i$ permanently plays caste. By Proposition 1 part 5, $i$’s outcomes are bounded in $S^{\prime}$, but we have already established that $i$’s outcomes are unbounded in $S$, so $S^{\prime}$ cannot be a strict Nash equilibrium. Then let $S^{\prime\prime}$ be any other joint strategy and let $i$ be the lowest-ranked caste player in $S^{\prime\prime}$ who is not lowest-ranked overall. By Proposition 1 part 3, $i$’s outcomes in $S^{\prime\prime}$ are permanently 0 and thus no better than they would be in any other joint strategy, including the strategy identical to $S^{\prime\prime}$ except that $i$ permanently takes turns. Therefore, $S^{\prime\prime}$ also cannot be a strict Nash equilibrium. ∎